OpenStudy (anonymous):

need help with this equation 5/(4 ) (3x-1/2)-2/(3 ) <7/3

OpenStudy (anonymous):

When you are solving and you have denominators, it is best to multiply every term by the least common denominator--- it will get rid of all denominators for you..

OpenStudy (anonymous):

Is the 2 under all of 3x - 1 or just under the 1?

OpenStudy (anonymous):

5 over 4 times 3x - 1/2 - 2/3 is less than 7 over 3

OpenStudy (anonymous):

So the least common denominator is 12 - so multiply every term by 12.

OpenStudy (anonymous):

@Blexting only under the 1

OpenStudy (anonymous):

looks like you have somone, good luck

OpenStudy (anonymous):

12(5/4) = 15 12 * 3x = 36x 12* (-1/2) = -6 12 * (-2/3) = -6 12 * (7/3) = 28

OpenStudy (anonymous):

Thx Schultzman8

OpenStudy (anonymous):

5/4 times what (3x - 1/2) in the original?

OpenStudy (anonymous):

ok i got 540x-90-6

OpenStudy (anonymous):

I am wrong... didn't get the parentheses correct

OpenStudy (anonymous):

yes @ 5/4 times what (3x - 1/2) in the original? ohhh ok

OpenStudy (anonymous):

god this class is killing me

OpenStudy (anonymous):

15/4x - 5/8 -2/3 < 7/3

OpenStudy (anonymous):

LCD = 24

OpenStudy (anonymous):

thats what I got I just did not like the end result

OpenStudy (anonymous):

24(15/4x) = 90x 24(-5/8) = -15 24(-2/3) = -16 24(7/3) = 56

OpenStudy (anonymous):

according to my text from the chapter the 7/3 is to be done after I get the first part done

OpenStudy (anonymous):

90x - 15 - 16 < 56 90x - 31 < 56 90x < 87 x < 87/90 (I don't like the answer either)

OpenStudy (anonymous):

so it should look like 90x-15-16 < 7/3

OpenStudy (anonymous):

If you multiply one of the terms by 24 you have to do it to all of them including the 7/3

OpenStudy (anonymous):

No.. that would not be correct.. it has to be done to everyterm. 90x - 15 - 16 < 56

OpenStudy (anonymous):

^he is right

OpenStudy (anonymous):

90x-31<56

OpenStudy (anonymous):

then times the the two denomanators 3x24 so I would have 90x - 15 - 16 < 72

OpenStudy (anonymous):

gosh I am confused LMAO

OpenStudy (anonymous):

90x<87 x would be a decimal

OpenStudy (anonymous):

my instructor is a douch

OpenStudy (anonymous):

so I got 0.966666

OpenStudy (anonymous):

Above you had 72... it was 24(7/3) so 3 goes into 24 8 times and then time 7 is 56

OpenStudy (anonymous):

remember, it is less than, not equal too

OpenStudy (anonymous):

therefore x= .966 or less

OpenStudy (anonymous):

Alright?

OpenStudy (anonymous):

so the solution set would be {x|x< .966}

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THX to both of you I have one more problem if u have time...

OpenStudy (anonymous):

alright, but can i get a medal?

OpenStudy (anonymous):

it is the elimination method

OpenStudy (anonymous):

of course how do i do that

OpenStudy (anonymous):

click good anwser for one of my anwsers

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

5x+5y= -11 7x-2y=19

OpenStudy (anonymous):

Now I will explain the Elimination method, to see if it helps you

OpenStudy (anonymous):

I get x=1.62

OpenStudy (anonymous):

I could not get Y cuzz i think x is wrong

OpenStudy (anonymous):

When adding two equations, you basically add all parts of them. Say you had the equations: 5x + 5y = -11 7x-2y= 19

OpenStudy (anonymous):

so 12x + 3y= 8

OpenStudy (anonymous):

I have not had any problems using it with other problems i just did not like the answer I got for X

OpenStudy (anonymous):

12x + 3y = 8.

OpenStudy (anonymous):

If you have those two equation - you have to eliminate one of the variables.. so multiply the first by 2 and the second by 5

OpenStudy (anonymous):

2(5x - 5y = 11) 5(7x - 2y = 19)

OpenStudy (anonymous):

elimination method is when you elou to cancel out on variable

OpenStudy (anonymous):

10x - 10y = 55 35x - 10y = 95

OpenStudy (anonymous):

Oops 10x + 10y = 55 35x - 10y = 95

OpenStudy (anonymous):

thats what I got blexting

OpenStudy (anonymous):

Add them up 45x = 150

OpenStudy (anonymous):

x = 3 1/3

OpenStudy (anonymous):

the 11 is a negative

OpenStudy (anonymous):

OOPS.. you are correct.

OpenStudy (anonymous):

10x + 10y = -55 35x - 10y = 95

OpenStudy (anonymous):

45x = 40 x = 40/45 x = 8/9

OpenStudy (anonymous):

i got 73/45

OpenStudy (anonymous):

ohhh ilike ur answer better then mine

OpenStudy (anonymous):

5x + 5y = -11 5(8/9) + 5y = -11

OpenStudy (anonymous):

40/9 + 5y = -11

OpenStudy (anonymous):

ok os not iplug 8/9 into

OpenStudy (anonymous):

4 4/9 + 5y = -11 5y = -15 4/9

OpenStudy (anonymous):

but i still have a fraction how does that solve

OpenStudy (anonymous):

I would do this instead.

OpenStudy (anonymous):

5(8/9) + 5y = -11 Multiply every term by 9... it gets rid of denominators 40 + 45y = -99

OpenStudy (anonymous):

45y = -139

OpenStudy (anonymous):

y = -3 4/45

OpenStudy (anonymous):

oh lord

OpenStudy (anonymous):

well thx for all ur help my assignment ti lto mor rowil tomorrow so I will keep all ur advice in mind I could not have figured it out hithout both of ur help!!!!!